# Automata Learning from Preference and Equivalence Queries

### v2

@@ Introduction
When this unification process is applied to a symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$, we are unifying $\mathcal{V}$ (the set of variables in the context $\Gamma$) according to the set of known equivalence constraints $\mathcal{C}_{EQ}$ and equivalence classes $\mathcal{C}_{EC}$ available in $\mathcal{C}$; and each variable in the table is replaced with its equivalence class representative (see Figure 2cdf after the large right curly brace). That is, for all $s$ in $(S\cup S\cdot\Sigma^{I})\cdot E$, we substitute $T(s)\leftarrow\mathbb{C}.\text{{rep}}$ if $T(s)\in\mathbb{C}$. In the resulting unified symbolic observation table, $T$ maps queried sequences to the set of representatives $\mathcal{R}$. Note $|\mathcal{R}|=|\mathcal{C}_{EC}|\leq|\Sigma^{O}|$.
@@
Besides constraints from preferences queries, the learner obtains constraints about the value of $T(c)$ from equivalence queries (Figure 2e to 2f) and adds them to $\mathcal{C}$. When first obtained, these constraints may possibly be expressed in terms of orphan variables, but during the process of unification, each orphan variable joins an equivalence class and is then replaced by its equivalence class representative in the constraint. Thus, after unification, all constraints in $\mathcal{C}$ are expressed in terms of equivalence class representatives.
@@
Definition 2.
A symbolic observation table is closed if $\textbf{{rows}}(S\cdot\Sigma^{I})\subseteq\textbf{{rows}}(S)$, where $\textbf{{rows}}(S)=\{\textbf{{row}}(s)|s\in S\}$, and $\textbf{{row}}(s)$ is the row in $\langle S,E,T;\mathcal{C},\Gamma\rangle$ indexed by $s$.
Definition 3.
A symbolic observation table is consistent if for all sequence pairs $s_{1}$ and $s_{2}$ where $\textbf{{row}}(s_{1})\equiv\textbf{{row}}(s_{2})$, all their transitions also remain equivalent with each other: $\textbf{{row}}(s_{1}\cdot\sigma)\equiv\textbf{{row}}(s_{2}\cdot\sigma)$ for all $\sigma\in\Sigma^{I}$.
Definition 4.
A symbolic observation table is unified if for all $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E$ it holds that $T(s\cdot e)\in\mathcal{R}$.
Making a Symbolic Hypothesis
If a symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is unified, closed, and consistent, a symbolic hypothesis can be made (Figure 2e). This construction is identical to $\textrm{L}^{*}$, except that the outputs are symbolic:
@@
Making a Concrete Hypothesis
The learner finds a satisfying solution $\Lambda$ to the set of constraints $\mathcal{C}$, while subject to the global constraint requiring the value of each representative to be in $\Sigma^{O}$ (Figure 2e). Thus, $\hat{L}$ becomes concrete.
@@
SymbolicFill  A symbolic fill (Algorithm 2 and Appendix Algorithm 4 and 5) (1) creates fresh variables for empty entries in the table, (2) asks preference queries, and (3) performs unification. If a sequence $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot\,E$ does not have an associated variable in the context $\Gamma$, then a fresh variable $\Gamma[s\cdot e]$ is created. Preference queries are executed to obtain the total ordering of $(S\cup(S\cdot\Sigma^{I}))\cdot E$. In our implementation, we use preference queries in place of comparisons in randomized quicksort and linear merge. Once every sequence in $(S\cup(S\cdot\Sigma^{I}))\cdot E$ has a variable, and once the preference queries have been completed, unification is performed. SymbolicFill is called on lines 2, 4 (in MakeClosedAndConsistent), and 7 (in ProcessCex.). Unification is shown in Appendix Algorithm 9.
@@
Ensuring Consistency   If the unified table is not consistent, then there exists a pair $s_{1},s_{2}\in S$ and $\sigma\in\Sigma^{I}$ for which $\textbf{{row}}(s_{1})\equiv\textbf{{row}}(s_{2})$ and $\textbf{{row}}(s_{1}\cdot\sigma)\not\equiv\textbf{{row}}(s_{2}\cdot\sigma)$, implying there is an $e\in E$ such that $T(s_{1}\cdot\sigma\cdot e)\not\equiv T(s_{2}\cdot\sigma\cdot e)$. To attempt to make the table consistent, add $\sigma\cdot e$ to $E$, and then perform a symbolic fill. Figure 2f shows inconsistency.
@@
Ensuring Closedness   If the unified table is not closed, then $\textbf{{rows}}(S\cdot\Sigma^{I})\not\subseteq\textbf{{rows}}(S)$. To attempt to make the table closed, find a row $\textbf{{row}}(s^{\prime})$ in $\textbf{{rows}}(S\cdot\Sigma^{I})$ but not in $\textbf{{rows}}(S)$. Add $s^{\prime}$ to $S$, update $S\cdot\Sigma^{I}$, then fill symbolically. Figure 2c to 2d shows a closure process.
@@
The closed and consistency checks occur in a loop (consistency first, closed second) inside MakeClosedAndConsistent until the table becomes unified, closed, and consistent. Then hypothesis $h=\langle\hat{Q},\hat{q}_{0},\Sigma^{I},\Sigma^{O},\hat{\delta},\hat{L}\rangle$ is generated by MakeHypothesis (Figure 2e) and is sent to the teacher via EquivalenceQuery (Figure 2e to 2f). If $h$ is wrong, then a counterexample $c$ is returned, as well as a new constraint on the value of $\hat{f}(c)$. The counterexample $c$ and all its prefixes are added to $S$, and then a symbolic fill is performed. Then the constraint on the value of $\hat{f}(c)$ is added to $\mathcal{C}$ as a constraint on the value of the representative at $T(c)$.
Theoretical Guarantees of Remap
We now cover the algorithmic guarantees of Remap when $\mathcal{T}$ uses exact equivalence queries, and show how sampling-based equivalence queries achieves PAC–identification.
@@
We first detail how Remap guarantees termination and yields a correct, minimal Moore machine that classifies sequence equivalently to $f$. If Remap terminates, then the final hypothesis must be correct, since termination occurs only if no counterexamples exist for the final hypothesis. Therefore, if the hypothesized Moore machine classifies all sequences correctly according to the teacher, it must be correct. Thus, proving termination implies correctness. See Appendix for sketches and proofs. Here, we assume the teacher provides feedback $\hat{f}(c)=f(c)$ with counterexample $c$.
Theorem 1.
@@
Theorem 3 (Query Complexity).
@@
Lemma 1, Theorem 2, and Corollary 1 imply Remap makes progress towards termination with every hypothesis, and termination occurs when specific conditions are satisfied; therefore its output must be correct. Theorem 3 indicates that REMAP learns the correct minimal automaton isomorphic to the target automaton in polynomial time.
Next, we show how Remap achieves PAC–identification when sampling-based equivalence queries are used.
Definition 5 (Probably Approximately Correct Identification).
@@
Theorem 5.
To achieve PAC-identification under Remap, given parameters $\epsilon$ and $d$, and if $f$ can be represented by a minimal Moore machine with $n$ states and $|\Sigma^{O}|$ classes, then teacher $\mathcal{T}$ needs to sample at least $\mathcal{O}\left(n+|\Sigma^{O}|+\frac{1}{\epsilon}\left((n+|\Sigma^{O}|)\ln\frac{1}{d}+(n+|\Sigma^{O}|)^{2}\right)\right)$ sequences i.i.d. from $\mathcal{D}$ over the entire run of Remap.
Theorem 4 and Theorem 5 imply Remap achieves PAC–identification for a choice of $\epsilon$ and $d$ as long as the teacher samples sufficient sequences per equivalence query.
Finally, Moore and Mealy machine equivalence permits converting the final hypothesis to a reward machine.
Learning Reward Machines from Preferences
@@
Markov Decision Processes  Decision making problems are often modeled by a Markov Decision Process (MDP), which is a tuple $\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$ where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the transition probability from state $s$ to $s^{\prime}$ via action $a$. The reward function $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ provides the associated scalar reward, and $0\leq\gamma\leq 1$ is a discount factor. In MDPs, the Markovian assumption is that transitions and rewards depend only upon the current state-action pair and the next state. However, Abel et al. (2021) show not all tasks are expressible using Markovian reward.
Non-Markovian Reward  Non-Markovian Reward Decision Processes (NMRDP) are identical to MDPs, except that $R:(\mathcal{S}\times\mathcal{A})^{*}\rightarrow\mathbb{R}$ is non-Markovian a reward function that depends on state-action history. This allows reward machines to model a class of non-Markovian reward functions.
@@
A single disjunctive normal formula (DNF) labeled transition can be used to summarize multiple transitions with identical $\Sigma^{O}$ labels (connecting a pair of states), since the elements of $2^{\mathcal{P}}$ are sets of propositions. Reward machines map label sequences to reward outputs and can be represented as $f:(2^{\mathcal{P}})^{*}\rightarrow\Sigma^{O}$, so Remap learns a RM by converting the output Moore machine to a reward machine.
Sequential Tasks in OfficeWorld and CraftWorld  Icarte et al. (2018) and Andreas et al. (2017) introduced the OfficeWorld and CraftWorld gridworld domains, respectively, and feature sequential tasks encoded as reward machines. OfficeWorld features 4 sequential tasks across several rooms with various objects available for an agent to interact with. Example tasks include (1) picking up coffee and mail and delivering them to a certain room, or (2) continuously patrolling between a set of rooms. CraftWorld is a 2D version of MineCraft, where the 10 sequential tasks involve the agent collecting materials and constructing tools or objects in a certain order while avoiding hazards.
Empirical Results
We evaluate the exact and PAC–identification (PAC-ID) versions of Remap and consider: first, how often is PAC-ID Remap correct? (Exact Remap is guaranteed to be correct). To answer this, we run experiments by applying PAC-ID Remap to learn reward machines (RMs), by converting the Moore machine into a RM. We measure empirical correctness with empirical probability of isomorphism and average regret. Second, how do exact and PAC-ID Remap scale? We measure preference query complexity as a function of the number of unique sequences queried, and present an example phase diagram of algorithm execution. We use Z3 (De Moura and Bjørner 2008) for the constraint solver.
Setup  We investigate these questions on 14 sequential tasks in the OfficeWorld and CraftWorld domains. The Appendix A contains domain specific details. We implement exact equivalence queries using a variant of the Hopcroft-Karp algorithm (Almeida, Moreira, and Reis 2009; Hopcroft and Karp 1971). We implement i.i.d. sequence sampling in sampling-based equivalence queries with the following process per sample: sample a length $L$ from a geometric distribution; then, construct an $L$-length sequence by drawing $L$ elements i.i.d. from a uniform distribution over $\Sigma^{I}$.
PAC–Identification Correctness Experiments

### v3

@@ 4 The Remap Algorithm
When this unification process is applied to a symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$, we are unifying $\mathcal{V}$ (the set of variables in the context $\Gamma$) according to the set of known equivalence constraints $\mathcal{C}_{EQ}$ and equivalence classes $\mathcal{C}_{EC}$ available in $\mathcal{C}$; and each variable in the table is replaced with its equivalence class representative (see Figure 3abde after the large right curly brace). That is, for all $s$ in $(S\cup S\cdot\Sigma^{I})\cdot E$, we substitute $T(s)\leftarrow\llbracket\mathbb{C}\rrbracket$ if $T(s)\in\mathbb{C}$. In the resulting unified symbolic observation table, $T$ maps queried sequences to the set of representatives $\mathcal{R}$. Note $|\mathcal{R}|=|\mathcal{C}_{EC}|\leq|\Sigma^{O}|$.
@@
Besides constraints from preferences queries, the learner obtains constraints about the value of $T(c)$ from equivalence queries (Figure 3c to 3d) and adds them to $\mathcal{C}$. When first obtained, these constraints may possibly be expressed in terms of orphan variables, but during the process of unification, each orphan variable joins an equivalence class and is then replaced by its equivalence class representative in the constraint. Thus, after unification, all constraints in $\mathcal{C}$ are expressed in terms of equivalence class representatives.
@@
Definition 7
Let $\textbf{{rows}}(S)=\{\textbf{{row}}(s)|s\in S\}$, where the row in $\langle S,E,T;\mathcal{C},\Gamma\rangle$ indexed by $s$ is $\textbf{{row}}(s)$. $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is closed if $\textbf{{rows}}(S\cdot\Sigma^{I})\subseteq\textbf{{rows}}(S)$.
Definition 8
$\langle S,E,T;\mathcal{C},\Gamma\rangle$ is consistent if for all sequence pairs $s_{1}$ and $s_{2}$ where $\textbf{{row}}(s_{1})\equiv\textbf{{row}}(s_{2})$, all their transitions also remain equivalent with each other: $\textbf{{row}}(s_{1}\cdot\sigma)\equiv\textbf{{row}}(s_{2}\cdot\sigma)$ for all $\sigma\in\Sigma^{I}$.
Definition 9
Table $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is unified if $T(s\cdot e)\in\mathcal{R}$ for all $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E$.
4.0.3 4.3 Making a Symbolic Hypothesis
If symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is unified, closed, and consistent, a symbolic hypothesis can be made (Figure 3c). This construction is identical to $\textrm{L}^{*}$, except that the outputs are symbolic:
@@
4.0.4 4.4 Making a Concrete Hypothesis
The learner finds a satisfying solution $\Lambda$ to the set of constraints $\mathcal{C}$, while subject to the global constraint requiring the value of each representative to be in $\Sigma^{O}$ (Figure 3c). Thus, $\hat{L}$ becomes concrete.
@@
SymbolicFill  A symbolic fill (Algorithm 2 and Appendix Algorithm 4 and 5) (i) creates fresh variables for empty entries in the table, (ii) asks preference queries, and (iii) performs unification. If a sequence $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot\,E$ does not have an associated variable in the context $\Gamma$, then a fresh variable $\Gamma[s\cdot e]$ is created. Preference queries are executed to obtain the total preordering of $(S\cup(S\cdot\Sigma^{I}))\cdot E$. In our implementation, we use preference queries in place of comparisons in randomized quicksort and linear merge. Once every sequence in $(S\cup(S\cdot\Sigma^{I}))\cdot E$ has a variable, and once the preference queries have been completed, unification is performed. SymbolicFill is called on lines 2, 4 (in MakeClosedAndConsistent), and 7 (in ProcessCex.). Unification is shown in Appendix Algorithm 9.
@@
Ensuring Consistency   If the unified table is not consistent, then there exists a pair $s_{1},s_{2}\in S$ and $\sigma\in\Sigma^{I}$ for which $\textbf{{row}}(s_{1})\equiv\textbf{{row}}(s_{2})$ and $\textbf{{row}}(s_{1}\cdot\sigma)\not\equiv\textbf{{row}}(s_{2}\cdot\sigma)$, implying there is an $e\in E$ such that $T(s_{1}\cdot\sigma\cdot e)\not\equiv T(s_{2}\cdot\sigma\cdot e)$. To attempt to make the table consistent, add $\sigma\cdot e$ to $E$, and then perform a symbolic fill. Figure 3d shows inconsistency, and Figure 3e shows a table made consistent through expansion of the suffix set.
@@
Ensuring Closedness   If the unified table is not closed, then $\textbf{{rows}}(S\cdot\Sigma^{I})\not\subseteq\textbf{{rows}}(S)$. To attempt to make the table closed, find a row $\textbf{{row}}(s^{\prime})$ in $\textbf{{rows}}(S\cdot\Sigma^{I})$ but not in $\textbf{{rows}}(S)$. Add $s^{\prime}$ to $S$, update $S\cdot\Sigma^{I}$, then fill symbolically. Figure 3a to 3b shows a closure process.
@@
The closed and consistency checks occur in a loop (consistency first, closed second) inside MakeClosedAndConsistent until the table becomes unified, closed, and consistent. Then hypothesis $h=\langle\hat{Q},\hat{q}_{0},\Sigma^{I},\Sigma^{O},\hat{\delta},\hat{L}\rangle$ is generated by MakeHypothesis (Figure 3c) and is sent to the teacher via EquivalenceQuery (Figure 3c to 3d). If $h$ is wrong, then a counterexample $c$ is returned, as well as feedback $\phi(c)$ which is interpreted as a new constraint on the value of $\hat{f}(c)$. The counterexample $c$ and all its prefixes are added to $S$, and then a symbolic fill is performed. Then the constraint on the value of $\hat{f}(c)$ is added to $\mathcal{C}$ as a constraint on the value of the representative at $T(c)$.
5 Theoretical Guarantees of Remap
@@
We now cover the algorithmic guarantees of Remap when $\mathcal{T}$ uses exact equivalence queries, and show how sampling-based equivalence queries achieves PAC–identification. We first detail how Remap guarantees termination and yields a correct, minimal Moore machine that classifies sequence equivalently to $f$. If Remap terminates, then the final hypothesis must be correct, since termination occurs only if no counterexamples exist for the final hypothesis. Therefore, if the hypothesized Moore machine classifies all sequences correctly according to the teacher, it must be correct. Thus, proving termination implies correctness. See Appendix for sketches and proofs. Here, we assume the teacher provides feedback $\hat{f}(c)=f(c)$ with counterexample $c$.
Theorem 5.1
@@
Proof
(Sketch) This sketch applies to the above lemma, theorem, and corollary about termination. Consider the sequence $\dots,h_{k-1},h_{k},\dots$ of hypotheses that Remap makes. For a given pair of consecutive hypotheses $(h_{k-1},h_{k})$, consider how the number of states $n$, and the number of known representative values $n_{\bullet}$ changes. Let $n^{*}$ be the number of states of the minimal Moore machine correctly classifying all sequences. Let $V^{*}\leq|\Sigma^{O}|$ be the upper bound on $|\mathcal{R}|$. Note that $0\leq n_{\bullet}\leq|\mathcal{R}|\leq V^{*}\leq|\Sigma^{O}|$ always holds. Through detailed case analysis on returned counterexamples, we can show that the change in $n_{\bullet}$, denoted by $\Delta n_{\bullet}$, must always be either 00 or $1$, and furthermore, if $\Delta n_{\bullet}=0$, then we must have $\Delta n\geq 1$. By the case analysis and tracking $n$ and $n_{\bullet}$, observe that if a counterexample $c$ is received from the teacher due to hypothesis $h$, then at least one of (a) $n<n^{*}$ or (b) $n_{\bullet}<V^{*}$ must be true. Since $\Delta n_{\bullet}$ and $\Delta n$ cannot simultaneously be 00, whenever a new hypothesis is made, progress must be made towards the upper bound of $(n^{*},V^{*})$. If the upper bound is reached, then the algorithm must terminate, since it is impossible to progress from the point $(n^{*},V^{*})$.∎
Theorem 5.3 (Query Complexity)
@@
Proof
$(n+m(n+|\Sigma^{O}|-1))(1+|\Sigma^{I}|)n=\mathcal{O}(mn^{2}).$
Lemma 1, Theorem 2, and Corollary 1 imply Remap makes progress towards termination with every hypothesis, and termination occurs when specific conditions are satisfied; therefore its output must be correct. Theorem 3 indicates that REMAP learns the correct minimal automaton isomorphic to the target automaton in polynomial time. Next, we show how Remap achieves PAC–identification when sampling-based equivalence queries are used.
Definition 10 (Probably Approximately Correct Identification)
@@
Proof
(Sketch) The probability $1-\epsilon_{k}$ of a sequence sampled from an arbitrary distribution $\mathcal{D}$ over $(\Sigma^{I})^{*}$ depends upon the distribution and the intersections of sets of sequences of the teacher and the learner’s $k$th hypothesis with the same classification values. The probability that the $k$th hypothesis misclassifies a sequence is $\epsilon_{k}$. If the teacher samples $m_{k}$ samples for the $k$th equivalence query, then an upper bound can be established for the case when $\epsilon_{k}\leq\epsilon$ for a given $\epsilon$. Since we know Remap executes at most $n+|\Sigma^{O}|-1$ equivalence queries, one can upper bound the probability that Remap terminates with an error by summing all probabilities of events that the teacher does not detect an error in at most $n+|\Sigma^{O}|-1$ equivalence queries. An exponential decaying upper bound can be found, and a lower bound for $m_{k}$ can be found in terms of $\epsilon,d,$ and $k$.∎
Theorem 5.5
To achieve PAC-identification under Remap, given parameters $\epsilon$ and $d$, and if $f$ can be represented by a minimal Moore machine with $n$ states and $|\Sigma^{O}|$ classes, then teacher $\mathcal{T}$ needs to sample at least
$\mathcal{O}\left(n+|\Sigma^{O}|+\frac{1}{\epsilon}\left((n+|\Sigma^{O}|)\ln\frac{1}{d}+(n+|\Sigma^{O}|)^{2}\right)\right)$
sequences i.i.d. from $\mathcal{D}$ over the entire run of Remap.
Proof
Since for the $k$th equivalence query, the teacher must sample at least $m_{k}\geq\left\lceil\frac{1}{\epsilon}(\ln\frac{1}{d}+k\ln 2)\right\rceil$ sequences in order to achieve PAC-identification, if the total number of samples is to be minimized while still achieving PAC-identification, then the teacher can just sample a quantity of sequences i.i.d. from $\mathcal{D}$ equal to the following total
$=n+|\Sigma^{O}|-1+\frac{1}{\epsilon}\left[(\ln\frac{1}{d})(n+|\Sigma^{O}|-1)+\ln 2\sum_{k=1}^{n+|\Sigma^{O}|-1}k\right]$
Theorem 4 and Theorem 5 imply Remap achieves PAC–identification for a choice of $\epsilon$ and $d$ as long as the teacher samples sufficient sequences per equivalence query. In particular, Theorem 5 indicates the total quantity of sequences sampled by the teacher to achieve PAC-identification depends on both $n$ the number of states, and $|\Sigma^{O}|$ the number of output classes. This contrasts with the result for PAC-identification of DFAs (Theorem 7 [5]), which depends only on $n$. Finally, since Remap outputs a Moore machine, one can leverage Moore and Mealy machine equivalence in order to convert the final hypothesis into a reward machine, as defined and covered in the next section.
6 Learning Reward Machines from Preferences
@@
Markov Decision Processes  Decision making problems are often modeled by a Markov Decision Process (MDP), which is a tuple $\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$ where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the transition probability from state $s$ to $s^{\prime}$ via action $a$. The reward function $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ provides the associated scalar reward, and $0\leq\gamma\leq 1$ is a discount factor. In MDPs, the Markovian assumption is that transitions and rewards depend only upon the current state-action pair and the next state. However, not all tasks are expressible using Markovian reward [2].
Non-Markovian Reward  Non-Markovian Reward Decision Processes are identical to MDPs, except that $R:(\mathcal{S}\times\mathcal{A})^{*}\rightarrow\mathbb{R}$ is non-Markovian a reward function that depends on state-action history. This allows reward machines to model a class of non-Markovian reward functions.
@@
A single disjunctive normal formula (DNF) labeled transition can summarize multiple transitions with identical $\Sigma^{O}$ labels (connecting a pair of states), since the elements of $2^{\mathcal{P}}$ are sets of propositions. Reward machines map label sequences to reward outputs and can be represented as $f:(2^{\mathcal{P}})^{*}\rightarrow\Sigma^{O}$, so Remap learns a reward machine by converting the output Moore machine to a reward machine.
Sequential Tasks in OfficeWorld and CraftWorld  Icarte et al. [25] and Andreas et al. [4] introduced the OfficeWorld and CraftWorld gridworld domains, respectively, and feature sequential tasks encoded as reward machines. OfficeWorld features 4 sequential tasks across several rooms with various objects available for an agent to interact with. Example tasks include (1) picking up coffee and mail and delivering them to a certain room, or (2) continuously patrolling between a set of rooms. CraftWorld is a 2D version of MineCraft, where the 10 sequential tasks involve the agent collecting materials and constructing tools or objects in a certain order while avoiding hazards.
7 Empirical Results
We evaluate the exact and PAC–identification (PAC-ID) versions of Remap and consider: first, how often is PAC-ID Remap correct? (Exact Remap is guaranteed to be correct). To answer this, we run experiments by applying PAC-ID Remap to learn reward machines (RMs), by converting the Moore machine into a RM. We measure empirical correctness with empirical probability of isomorphism and average regret. Second, how do exact and PAC-ID Remap scale? We measure preference query complexity as a function of the number of unique sequences queried, and present an example phase diagram of algorithm execution. We use Z3 [15] for the constraint solver.
Setup  We investigate these questions on 14 sequential tasks in the OfficeWorld and CraftWorld domains. The Appendix 0.A.1 contains domain specific details. We implement exact equivalence queries (Definition 5) using a variant of the Hopcroft-Karp algorithm [3, 24]. We implement i.i.d. sequence sampling in sampling-based equivalence queries with the following process per sample: sample a length $L$ from a geometric distribution; then, construct an $L$-length sequence by drawing $L$ elements i.i.d. from a uniform distribution over $\Sigma^{I}$. In both the exact and sampling-based equivalence queries, strong feedback is used.
7.1 PAC–Identification Correctness Experiments